OpenStudy (anonymous):

The region in the first quadrant bounded by the graph of y=sec(x), x=3.14(pi)/4, and the axes is rotated about the x-axis. What is the volume of the solid generated?

OpenStudy (accessdenied):

We'll use a method called the "disk" method. We're basically taking the sum of the volumes of very thin disks over the interval \([0,\frac{\pi}{4}]\). These disks have a radius of \( f(x) \) and a width of \( dx \). The equation for the volume of a single disk at a given \( x \) is \( \color{red}V = \pi (f(x))^2 dx \).  So, the sum of all of these thin disks is the definite integral on the interval \([0,\frac{\pi}{4}]\): \[~ \int_{0}^{\frac{\pi}{4}} \color{red}V = \int_{0}^{\frac{\pi}{4}} \pi (f(x))^2 dx \] Here, \(f(x) = sec(x)\). \[~ \int_{0}^{\frac{\pi}{4}} \pi sec^2(x) dx = \pi \int_{0}^{\frac{\pi}{4}} sec^2(x) dx \] You can bring out the \( \pi \) since it is just a constant.